A farmer has a triangular field besides 120 yards, 170 yards, and 220 yards. Find the area of the field in square yards.
Answer by addingup(3677)

Knowing all three sides we can use Hero's formula:
Let a, b, c be the sides
Let p = 1/2 perimeter
``````````````````````````````````````````````
Perimeter: a+b+c = 120+170+220 = 510
p = 510/2 = 255
`````````````````````````````````````````````
Hero's formula:
√p(p-a)(p-b)(p-c) = √255(255-120)(255-170)(255-220) = 10,120 or if you want to be super precise: 10,119.99876
John